Briana uses the following equation to model the number of infected people each week.
Y=50 e^0.3t. Y represents the number of infected people and t represents time in weeks. In how many weeks will the number of infected people reach 2500.

Answer by Cromlix(4381)

Hi there,
Y=50 e^0.3t.
In how many weeks will the number
of infected people reach 2500?
50 e^0.3t = 2,500
e^0.3t = 2,500/50
e^0.3t = 50
0.3t loge e = loge(50)
(loge e = 1)
0.3t = loge (50)
t = loge (50)/0.3
t = 13.04 weeks
Hope this helps :-)